# Unleashing the Power of CNN and Transformer for Balanced RGB-Event Video Recognition

### v1

@@ Abstract
I Introduction
Event-based vision has drawn more and more attention in recent years [1, 2, 3] due to its significant advantages on high dynamic range, low energy consumption, spatial sparse and temporal dense. It has been introduced in many computer vision tasks, such as object recognition [4, 5], detection and tracking [2, 3, 6], and image enhancement and reconstruction [7, 8, 9, 10, 11]. Different from standard RGB sensors which record the light intensity for all pixels in a global exposure way (i.e., synchronously), Event cameras (also called Dynamic Vision Sensors, DVS) emit a spike/event only when the variation of light intensity in a pixel exceeds a given threshold. In other words, the Event cameras asynchronously record the scene using spike/event points similar to point clouds. A comparison between the RGB frames and Event streams is provided in Fig. 2.
Although the Event camera performs well on the aforementioned aspects, however, it captures fewer signals when the target object is motionless or static and also fails to record the color or detailed texture information. These issues make it hard to utilize only Event cameras for some practical scenarios. Hence, some researchers attempt to combine the RGB and Event cameras for a more robust and strong performance [12, 3, 13]. In this work, we focus on the task of RGB-Event based pattern recognition which is a fundamental research problem for the Event-based vision. The key challenge is how to encode and fuse these two heterogeneous modalities simultaneously for a better tradeoff between the accuracy and model parameters.
@@ I Introduction
According to our observation, we find that previous RGB-Event based recognition algorithms usually process the dual modalities using CNN (Convolutional Neural Networks), SNN (Spiking Neural Networks), GNN (Graph Neural Networks), or Transformer Networks, due to the diverse representations of Event streams. To be specific, Wang et al. [12] propose the SSTFormer which encodes the RGB and Event data using Memory-Support Transformer and SNN and fuse them using a bottleneck mechanism. Although a good performance can be obtained, however, their model contains more than 600M learnable parameters which make it difficult to optimize during actual deployment. For the models with less than 100M parameters like TSM [14], TAM [15], and Video SwinTransformer [16], as noted in Fig. 1, we can find that their results are relatively poor on the public PokerEvent [12] benchmark dataset. These observations inspired us to think about how to achieve a good balance between the model parameters and recognition performance for the RGB-Event based recognition?
@@
The rest of this paper is organized as follows: In section II, we give a brief introduction to the most related works. Then, we will introduce our proposed TSCFormer framework in section III with a focus on the overview, input representation, temporal shift CNN sub-network, BridgeFormer sub-network, classification head, and loss function. We conduct extensive experiments on two datasets which will be introduced in section IV-A. The implementation details, comparisons with SOTA models, ablation studies, parameter analysis, visualizations, and limitation analysis are all described in this part. Finally, we conclude this paper and propose possible research directions for the next steps in section V.
II Related Works
@@
In this section, we give a brief review of Event-based Recognition, RGB-Event based Recognition, and Transformer Networks. More related works can be found in the following surveys [1, 17, 18] and paper list  [^1] https://github.com/Event-AHU/Event_Camera_in_Top_Conference.
II-A Event-based Recognition
@@
In recent years, Event-based technology has garnered widespread attention in the fields of computer vision and machine learning. Significant progress has been made in Event-based recognition, providing valuable experience and theoretical foundations for the development of Event-based recognition technology. To be specific, Cho et al. [19] propose a joint learning framework for Event-based object recognition and image reconstruction, which conducts joint learning without requiring paired images and labels. Innocenti et al. [20] propose a concise representation of event data called Temporal Binary Representation, which leverages the conversion of binary event sequences into frames to encode both spatial and temporal information. Their approach enables the adjustment of information loss and memory footprint, rendering it suitable for real-time applications. The authors of [21] propose a test time adaptation algorithm for event-based object recognition, referred to as Ev-TTA. It enables the classifier to adapt to new, unseen, and widely varying environments during the testing phase. Cannici et al. [22] propose an attention mechanism algorithm that monitors event activity in a scene and extracts patches from reconstructed frames, for image classification, capable of recognizing objects within the reconstructed frames. Wang et al. [23] propose an Event-based gait recognition method, Ev-Gait. This algorithm effectively eliminates noise from the event stream by enhancing motion consistency and utilizes deep neural networks to recognize gait from asynchronous and sparse event data. Liu et al. [24] propose a hierarchical SNN architecture for Event-based action recognition using motion information. Compared with these works, our proposed TSCFormer is a multi-modal framework that models the RGB and Event data simultaneously using temporal shift based CNN and BridgeFormer module.
II-B RGB-Event based Recognition
@@
Multi-modal based research has been widely exploited in many downstream tasks [25, 26, 27]. By integrating RGB images and Event data, RGB-Event based recognition can provide a more robust recognition performance. In comparison to traditional Event-based recognition, RGB-Event recognition places an emphasis on the fusion of visual information and Event data to achieve a comprehensive and enriched understanding of scenes and actions. Previous work provides an empirical and theoretical basis for the development of Event-based recognition techniques. Specifically, Huang et al. [28] propose to utilize a cross-attention module on the RGB-Event bimodal data to endow the network with cross-modal feature selection capabilities, while the self-attention module combines nearby local features. Wang et al. [12] propose an RGB frame-event recognition framework for the simultaneous fusion of RGB frames and event streams, termed SSTFormer. It utilizes a memory-supported Transformer network for RGB frame encoding and a spiking neural network for raw event stream encoding. Li et al. [4] introduce a Semantic-Aware Frame-Event fusion based pattern recognition framework, called SAFE, based on a large-scale vision-language pre-trained model. This model integrates RGB frames, Event data, and semantic label sets into a unified framework. Compared to previous works, the TSCFormer proposed in this work processes the RGB and Event data using temporal shift CNN and further enhances the local CNN features using the BridgeFormer module. Note that, the BridgeFormer proposed for the global feature learning is lightweight and achieves a better tradeoff between the accuracy and model complexity.
II-C Transformer Network
@@
In recent years, the Transformer [29] network is increasingly applied to fuse information from different modalities [30, 31, 32, 33], effectively processing and integrating diverse types of data. This demonstrates the potential of the Transformer architecture in handling multi-modal data. As a result, there has been growing interest in exploring and advancing multi-modal Transformer networks for a wide range of applications. Peng et al. [34] propose a hybrid neural network that combines the CNN and Transformer for RGB-based representation learning, termed ConvFormer. It encodes the local and global features by inputting the RGB image into the CNN and Transformer, respectively. Bottleneck fusion based multi-modal framework is proposed in [35] by Arsha et al. which achieves a better balance between the performance and computational cost. Zhang et al. [36] explore RGB-X semantic segmentation for the first time using five multi-modal sensor data combinations: RGB-Depth, RGB-Thermal, RGB-Polarization, RGB-Event, and RGB-LiDAR. The authors propose an RGB-X semantic segmentation framework called CMX, which features cross-modal feature calibration and fusion modules, interleaved cross-attention, and mixed-channel embedding to enhance global reasoning. Zhao et al. [37] design a Transformer-based Fusion Module (TFM) to explore the relationship between the spectral representations of hyper-spectral and RGB data within this framework. The author of [38] employs the Transformer network to learn the fusion of time-domain multi-view features to obtain a hierarchical structure in a coarse to fine manner. Wang et al. [39] introduce a transformer-based RGB-D saliency detection framework, which globally integrates cross-modal and scale features simultaneously. Zhou et al. [40] propose a multi-spectral fusion transformer network to highlight the intra-spectral correlation and inter-spectral complementary of RGB-T. Li et al. [41] propose a feature fusion block that pays mutual attention to learning a joint feature representation for classification. The authors of [42] propose a multi-modal token fusion method (TokenFusion) that effectively integrates multiple modalities, allowing the transformer to learn correlations. Prakash et al. [43] propose a Multi-Modal Fusion Transformer (TransFuser) to integrate the global context of the 3D scene into the feature extraction layers of various modalities. Inspired by these works, in this paper, we propose a novel framework that facilitates the information propagation between local convolutional RGB/Event features and global Transformer representations in a relatively lightweight way.
@@
Figure 3: An overview of our proposed RGB-Event based pattern recognition framework, termed TSCFormer. Specifically, we first adopt a stem network to process the RGB frames and event streams into feature embeddings. Then, the local features of RGB and Event data can be obtained using temporal-shift-based CNN (ResNet50 [44] is used in this work). To get the global representations, we randomly initialize a set of global tokens and feed them into the BridgeFormer module. Therefore, the output global tokens will be fed into RGB and Event branch for fusion, and the subsequent BridgeFormer module for further information aggregation. Finally, we fuse all three branches and feed them into a fully connected layer for classification.
III Our Proposed Method
@@
III-A Overview
As shown in Fig. 3, given the RGB frames and Event streams, we first adopt a stem net which contains a convolutional layer and max-pooling layer to obtain the feature embeddings. Then, the two feature maps are fed into temporal-shift-based CNN modules to get the local feature representations. Meanwhile, a set of global tokens is randomly initialized and fed into the BridgeFormer layer. Note that, the local RGB/Event feature maps are concatenated and transformed into key and value tokens, and the global tokens are transformed into query tokens. Then, a cross-attention layer is adopted to fuse the local and global features and a Transformer layer is used to learn the global tokens. The output features will be concatenated with RGB and Event CNN representations after feature transformation via a fully connected layer and reshape operation. After that, we adopt a convolutional layer with kernel size $1\times 1$ to fuse these features and feed them into the subsequent temporal shift CNN blocks for a similar operation. Finally, we flatten the features of the two modalities and map them into category labels using the classifier. More detailed computing procedures will be introduced in the following subsections.
III-B Input Representation
@@
III-C Temporal Shift CNN Sub-Network
After we get the embedded feature maps of RGB and Event frames, i.e., $F_{I}$ and $F_{E}$, we feed them into temporal-shift-based CNN modules, respectively. As noted in TSM [14], this simple operation can achieve the performance of 3D CNN but maintain 2D CNN’s complexity. Specifically speaking, the temporal-shift module will shift part of the channels along the temporal dimension, therefore, facilitating information exchange between neighboring frames. An illustration of the temporal shift module is provided in Fig. 4. Then, the shifted features $\hat{F}_{I}$ and $\hat{F}_{E}$ will be fed into the CNN blocks (ResNet50 [44] is adopted in this work) for enhanced feature extraction. In our practical implementation, four residual blocks are adopted for the feature extraction. For each block $i\in\{1,2,3,4\}$, we concatenate the RGB and Event features [$\hat{F}_{I}^{i}$, $\hat{F}_{E}^{i}$] along the channel dimension and feed into the BridgeFormer module which will be introduced in the next subsection.
@@ III-C Temporal Shift CNN Sub-Network
III-D BridgeFormer Sub-Network
The aforementioned temporal-shift CNN block learns the local features well, however, the global long-range relations can be better obtained using the self-attention-based Transformer network as validated in many previous works [34][43]. In this work, we propose a novel BridgeFormer that efficiently fuses and enhances the RGB and Event features. More specifically, we randomly initialize a set of global tokens $\mathcal{Z}\in\mathbb{R}^{B\times L\times D}=\{Z_{1},Z_{2},...,Z_{B}\}$ and treat it as the query feature $\mathcal{Q}$. The concatenated RGB-Event features [$\hat{F}_{I}^{i}$, $\hat{F}_{E}^{i}$] are transformed into key and value features, i.e., $\mathcal{K}$ and $\mathcal{V}$. Then, a cross-attention layer is utilized to fuse these tokens and the detailed procedure can be formulated as:
@@ III-D BridgeFormer Sub-Network
We also introduce the Feed-Forward Network (FFN) to get the enhanced feature tokens inspired by ViT [45] and Transformer [29]. The output will be fused with RGB and Event CNN features, respectively, and also passed into the next BridgeFormer module for further processing.
@@
III-E Classification Head & Loss Function

### v2

@@ Abstract
1 Introduction
Event-based vision has drawn more and more attention in recent years gallego2020eventsurvey ; wang2023eventvot ; wang2023visevent due to its significant advantages on high dynamic range, low energy consumption, spatial sparse and temporal dense. It has been introduced in many computer vision tasks, such as object recognition li2023semantic ; wang2022hardvs , detection and tracking wang2023eventvot ; wang2023visevent ; gehrig2023recurrentVFormer , and image enhancement and reconstruction teng2022nest ; zhu2022eventreconstruction ; jiang2023eventLIenhance ; ding2023emlb ; wu2020probabilistic . Different from standard RGB sensors which record the light intensity for all pixels in a global exposure way (i.e., synchronously), Event cameras (also called Dynamic Vision Sensors, DVS) emit a spike/event only when the variation of light intensity in a pixel exceeds a given threshold. In other words, the Event cameras asynchronously record the scene using spike/event points similar to point clouds.
When the light intensity change at a pixel exceeds the set threshold, the camera records an event. Each event is recorded as a quadruple (x, y, t, p), where x and y represent the spatial coordinates where the event occurs, t is the timestamp, and p represents the polarity of the event. Generally, +1 indicates an increase in brightness, and -1 indicates a decrease in brightness. This characteristic enables the event camera to asynchronously output an event stream related to the change in light intensity and to respond to the dynamic changes in the scene within an extremely short period of time. Event cameras have a high temporal resolution and can focus on the changes in the image on a microsecond time scale. In scenarios where objects are moving at high speed or there are rapid changes in illumination, event cameras are able to capture the changes in objects and record the motion information. Compared with each frame of images recorded by traditional cameras, event cameras only output the pixel information of pixels whose brightness has changed. This enables event cameras to require less bandwidth. A comparison between the RGB frames and Event streams is provided in Fig. 2.
Although the Event camera performs well on the aforementioned aspects, it captures fewer signals when the target object is motionless or static. And only a few event cameras are capable of recording color or detailed texture information, such as the DAVIS346 Scheerlinck2019CED: .
These issues make it hard to utilize only Event cameras for some practical scenarios. Hence, some researchers attempt to combine the RGB and Event cameras for a more robust and strong performance wang2023sstformer ; wang2023visevent ; jing2021VSREvent . In this work, we focus on the task of RGB-Event based pattern recognition which is a fundamental research problem for the Event-based vision. The key challenge is how to encode and fuse these two heterogeneous modalities simultaneously for a better tradeoff between the accuracy and model parameters.
@@ 1 Introduction
According to our observation, we find that previous RGB-Event based recognition algorithms usually process the dual modalities using CNN (Convolutional Neural Networks), SNN (Spiking Neural Networks), GNN (Graph Neural Networks), or Transformer Networks, due to the diverse representations of Event streams. To be specific, Wang et al. wang2023sstformer propose the SSTFormer which encodes the RGB and Event data using Memory-Support Transformer and SNN and fuse them using a bottleneck mechanism. Although a good performance can be obtained, however, their model contains more than 600M learnable parameters which make it difficult to optimize during actual deployment. For the models with less than 100M parameters like TSM lin2019tsm , TAM liu2021tam , and Video SwinTransformer liu2022video , as noted in Fig. 1, we can find that their results are relatively poor on the public PokerEvent wang2023sstformer benchmark dataset. These observations inspired us to think about how to achieve a good balance between the model parameters and recognition performance for the RGB-Event based recognition?
@@
The rest of this paper is organized as follows: In section 2, we give a brief introduction to the most related works. Then, we will introduce our proposed TSCFormer framework in section 3 with a focus on the overview, input representation, temporal shift CNN sub-network, BridgeFormer sub-network, classification head, and loss function. We conduct extensive experiments on two datasets which will be introduced in section 4.1. The implementation details, comparisons with SOTA models, ablation studies, parameter analysis, visualizations, and limitation analysis are all described in this part. Finally, we conclude this paper and propose possible research directions for the next steps in section 5.
2 Related Works
@@
In this section, we give a brief review of Event-based Recognition, RGB-Event based Recognition, and Transformer Networks. More related works can be found in the following surveys gallego2020eventsurvey ; han2022transformersurvey ; wang2023MMPTMs and paper list  [^1] https://github.com/Event-AHU/Event_Camera_in_Top_Conference.
2.1 Event-based Recognition
@@
In recent years, Event-based technology has garnered widespread attention in the fields of computer vision and machine learning. Significant progress has been made in Event-based recognition, providing valuable experience and theoretical foundations for the development of Event-based recognition technology. To be specific, Cho et al. cho2023label propose a joint learning framework for Event-based object recognition and image reconstruction, which conducts joint learning without requiring paired images and labels. Innocenti et al. innocenti2021temporal propose a concise representation of event data called Temporal Binary Representation, which leverages the conversion of binary event sequences into frames to encode both spatial and temporal information. Their approach enables the adjustment of information loss and memory footprint, rendering it suitable for real-time applications. The authors of Kim_2022_CVPR propose a test time adaptation algorithm for event-based object recognition, referred to as Ev-TTA. It enables the classifier to adapt to new, unseen, and widely varying environments during the testing phase. Cannici et al. cannici2019attention propose an attention mechanism algorithm that monitors event activity in a scene and extracts patches from reconstructed frames, for image classification, capable of recognizing objects within the reconstructed frames. Wang et al. Wang_2019_CVPR propose an Event-based gait recognition method, Ev-Gait. This algorithm effectively eliminates noise from the event stream by enhancing motion consistency and utilizes deep neural networks to recognize gait from asynchronous and sparse event data. Liu et al. liu2021event propose a hierarchical SNN architecture for Event-based action recognition using motion information. Compared with these works, our proposed TSCFormer is a multi-modal framework that models the RGB and Event data simultaneously using temporal shift based CNN and BridgeFormer module.
2.2 RGB-Event based Recognition
@@
Multi-modal based research has been widely exploited in many downstream tasks islam2022maven ; mai2022multimodalbottleneck ; wang2022mfgnet . By integrating RGB images and Event data, RGB-Event based recognition can provide a more robust recognition performance. In comparison to traditional Event-based recognition, RGB-Event recognition places an emphasis on the fusion of visual information and Event data to achieve a comprehensive and enriched understanding of scenes and actions. Previous work provides an empirical and theoretical basis for the development of Event-based recognition techniques. Specifically, Huang et al. huang2022vefnet propose to utilize a cross-attention module on the RGB-Event bimodal data to endow the network with cross-modal feature selection capabilities, while the self-attention module combines nearby local features. Wang et al. wang2023sstformer propose an RGB frame-event recognition framework for the simultaneous fusion of RGB frames and event streams, termed SSTFormer. It utilizes a memory-supported Transformer network for RGB frame encoding and a spiking neural network for raw event stream encoding. Li et al. li2023semantic introduce a Semantic-Aware Frame-Event fusion based pattern recognition framework, called SAFE, based on a large-scale vision-language pre-trained model. This model integrates RGB frames, Event data, and semantic label sets into a unified framework. Compared to previous works, the TSCFormer proposed in this work processes the RGB and Event data using temporal shift CNN and further enhances the local CNN features using the BridgeFormer module. Note that, the BridgeFormer proposed for the global feature learning is lightweight and achieves a better tradeoff between the accuracy and model complexity.
2.3 Transformer Network
@@
In recent years, the Transformer vaswani2017attention network is increasingly applied to fuse information from different modalities Truong_2021_ICCV ; radford2021learning ; He_2022_CVPR ; Neimark_2021_ICCV , effectively processing and integrating diverse types of data. This demonstrates the potential of the Transformer architecture in handling multi-modal data. As a result, there has been growing interest in exploring and advancing multi-modal Transformer networks for a wide range of applications. Peng et al. peng2021conformer propose a hybrid neural network that combines the CNN and Transformer for RGB-based representation learning, termed ConvFormer. It encodes the local and global features by inputting the RGB image into the CNN and Transformer, respectively. Bottleneck fusion based multi-modal framework is proposed in nagrani2021AttBottleNeck by Arsha et al. which achieves a better balance between the performance and computational cost. Zhang et al. zhang2023cmx explore RGB-X semantic segmentation for the first time using five multi-modal sensor data combinations: RGB-Depth, RGB-Thermal, RGB-Polarization, RGB-Event, and RGB-LiDAR. The authors propose an RGB-X semantic segmentation framework called CMX, which features cross-modal feature calibration and fusion modules, interleaved cross-attention, and mixed-channel embedding to enhance global reasoning. Zhao et al. zhao2022tftn design a Transformer-based Fusion Module (TFM) to explore the relationship between the spectral representations of hyper-spectral and RGB data within this framework. The author of bozic2021transformerfusion employs the Transformer network to learn the fusion of time-domain multi-view features to obtain a hierarchical structure in a coarse to fine manner. Wang et al. wang2021transformer introduce a transformer-based RGB-D saliency detection framework, which globally integrates cross-modal and scale features simultaneously. Zhou et al. zhou2022multispectral propose a multi-spectral fusion transformer network to highlight the intra-spectral correlation and inter-spectral complementary of RGB-T. Li et al. li2021trear propose a feature fusion block that pays mutual attention to learning a joint feature representation for classification. The authors of Wang_2022_CVPR propose a multi-modal token fusion method (TokenFusion) that effectively integrates multiple modalities, allowing the transformer to learn correlations. Prakash et al. Prakash_2021_CVPR propose a Multi-Modal Fusion Transformer (TransFuser) to integrate the global context of the 3D scene into the feature extraction layers of various modalities. Inspired by these works, in this paper, we propose a novel framework that facilitates the information propagation between local convolutional RGB/Event features and global Transformer representations in a relatively lightweight way.
@@
Figure 3: An overview of our proposed RGB-Event based pattern recognition framework, termed TSCFormer. Specifically, we first adopt a stem network to process the RGB frames and event streams into feature embeddings. Then, the local features of RGB and Event data can be obtained using temporal-shift-based CNN (ResNet50 He_2016_CVPR is used in this work). To get the global representations, we randomly initialize a set of global tokens and feed them into the BridgeFormer module. Therefore, the output global tokens will be fed into RGB and Event branch for fusion, and the subsequent BridgeFormer module for further information aggregation. Finally, we fuse all three branches and feed them into a fully connected layer for classification.
3 Our Proposed Method
@@
3.1 Overview
As shown in Fig. 3, given the RGB frames and Event streams, we first adopt a stem net which contains a convolutional layer and max-pooling layer to obtain the feature embeddings. Then, the two feature maps are fed into temporal-shift-based CNN modules to get the local feature representations. Meanwhile, a set of global tokens is randomly initialized and fed into the BridgeFormer layer. Note that, the local RGB/Event feature maps are concatenated and transformed into key and value tokens, and the global tokens are transformed into query tokens. Then, a cross-attention layer is adopted to fuse the local and global features and a Transformer layer is used to learn the global tokens. The output features will be concatenated with RGB and Event CNN representations after feature transformation via a fully connected layer and reshape operation. After that, we adopt a convolutional layer with kernel size $1\times 1$ to fuse these features and feed them into the subsequent temporal shift CNN blocks for a similar operation. Finally, we flatten the features of the two modalities and map them into category labels using the classifier. Fig. 5 shows the specific details of the network architecture. More detailed computing procedures will be introduced in the following subsections.
3.2 Input Representation
@@
3.3 Temporal Shift CNN Sub-Network
After we get the embedded feature maps of RGB and Event frames, i.e., $F_{I}$ and $F_{E}$, we feed them into temporal-shift-based CNN modules, respectively. As noted in TSM lin2019tsm , this simple operation can achieve the performance of 3D CNN but maintain 2D CNN’s complexity. Specifically speaking, the temporal-shift module will shift part of the channels along the temporal dimension, therefore, facilitating information exchange between neighboring frames. An illustration of the temporal shift module is provided in Fig. 4. Then, the shifted features $\hat{F}_{I}$ and $\hat{F}_{E}$ will be fed into the CNN blocks (ResNet50 He_2016_CVPR is adopted in this work) for enhanced feature extraction. In our practical implementation, four residual blocks are adopted for the feature extraction. For each block $i\in\{1,2,3,4\}$, we concatenate the RGB and Event features [$\hat{F}_{I}^{i}$, $\hat{F}_{E}^{i}$] along the channel dimension and feed into the BridgeFormer module which will be introduced in the next subsection.
@@ 3.3 Temporal Shift CNN Sub-Network
3.4 BridgeFormer Sub-Network
The aforementioned temporal-shift CNN block learns the local features well, however, the global long-range relations can be better obtained using the self-attention-based Transformer network as validated in many previous works peng2021conformer Prakash_2021_CVPR . In this work, we propose a novel BridgeFormer that efficiently fuses and enhances the RGB and Event features. More specifically, we randomly initialize a set of global tokens $\mathcal{Z}\in\mathbb{R}^{B\times L\times D}=\{Z_{1},Z_{2},...,Z_{B}\}$ and treat it as the query feature $\mathcal{Q}$. The concatenated RGB-Event features [$\hat{F}_{I}^{i}$, $\hat{F}_{E}^{i}$] are transformed into key and value features, i.e., $\mathcal{K}$ and $\mathcal{V}$. Then, a cross-attention layer is utilized to fuse these tokens and the detailed procedure can be formulated as:
@@ 3.4 BridgeFormer Sub-Network
We also introduce the Feed-Forward Network (FFN) to get the enhanced feature tokens inspired by ViT dosovitskiy2020image and Transformer vaswani2017attention . The output will be fused with RGB and Event CNN features, respectively, and also passed into the next BridgeFormer module for further processing.
@@
3.5 Classification Head & Loss Function